A sample of hard water contains 96 ppm of SO4 2- and 183 pppm of HCO3- and Ca 2+ as the only cation.How many moles of CaO will be required to remove HCO3-from 1000 kg of this water?If this water is treated with the amount of CaO calculated above,what will be the concentration (inppm) of residual Ca 2+ ions in one litre of the treated water are completely exchanged with hydrogen ions,what will be the pH ?

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

In 106 g(= 1000 kg) of the given hard water, we have amount of SO4 2– ions = 96 g
amount of HCO3 – ions = 183 g
So amount of SO4 2– ions = 96 g mol  = 1 mol
and amount of HCO3 – ions = 61 g mol
 = 3 mol
These ions are present as CaSO4 and Ca(HCO3)2.
Hence, amount of Ca2+ ions =  2.5 mol
The addition of CaO causes the following reactions:
CaO + Ca(HCO3)2 → 2CaCO3 + H2O
1.5 mol of CaO will be required for the removal of 1.5 mol of Ca(HCO3)2 in form of CaCO3. In the treated water, only CaSO4 is present now.
Thus, 1 mol of Ca2+ ions will be present in 106 g of water. Hence, its concentration will be 40 ppm.

Molarity of Ca2+ ions in the treated water will be 10–3 mol l–1. If the Ca2+ ions are exchanged by H+ ions then,
Molartiy of H+ in the treated water = 2 × 10–3 M
Thus, pH = – log(2 × 10–3) = 2.7

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi